A sports car must achieve a velocity of 100m/s after passing the entry tollgate of an expressway. If it can do this after traveling 1km, what must it's acceleration ve?


1
Expert's answer
2021-11-07T19:25:26-0500

"v_0=0"

"s _1= 1km=1000\frac{m}{s}"

"v_1= 100\frac{m}{s}"

"v_1 = v_0+at=at"

"at = 100"

"s = v_0t+\frac{at^2}{2}"

"s_1=\frac{at^2}{2}= at*\frac{t}{2}=50t"

"t = 20s"

"a = \frac{v_1}{t}=\frac{100}{20}=5\frac{m}{s^2}"


"\text{Answer: }5\frac{m}{s^2}"